# version 0.6

### v3

@@ 10.1 Nonstochastic System Identification
For $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, assuming that
@@
    the system is $(k,\kappa)$-strongly controllable system,
@@
    the perturbations $\mathbf{w}_{t}$ are chosen obliviously (i.e. do not depend on $\mathbf{u}_{t}$),
Algorithm 5 outputs a pair $(\widehat{A},\widehat{B})$ that satisfies, with probability $1-\Delta$,
@@
First note that for any $t$, $\mathbf{x}_{t}$ is bounded as
$\|\mathbf{x}_{t}\|=\left\|\sum_{i>1}A^{i-1}\left(\mathbf{w}_{t-i}+B\eta_{t-i}\right)\right\|\leq 2\sqrt{d}\gamma.$
@@
By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 10.4.
@@
Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. Hence, we will invoke the above stated lemma on the normal equations, i.e. $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$. First, observe that for any $\varepsilon<1$
@@ 10.2 From Indentification to Nonstochastic Control
Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well do GPC and GRC algorithms perform on the original system $(A,B)$ when provided with only approximate estimates of the same? We briefly state a result to this extent, whose proof we leave to the explorations of an interested reader.
@@
This result translates to a regret bound of $\tilde{O}(T^{2/3})$ for nonstochastic control of unknown, yet stable, linear systems, by first running the nonstochastic system identification procedure, and subsequently using the estimates thus obtained to run GPC. (Details left as Exercise 10.4
A similar statement holds for the GRC control algorithm, which in addition does not require strong controllability, since it can operate with estimates of the Markov operators alone (using Lemma 9.3), instead of estimates of the system matrices.
@@ 10.4 Exercises
10.2. Design an analogous control procedure as that for Exercise 10.4 with an identical regret upper bound (upto constants), when the horizon $T$ is not known in advance.
@@
Chapter 11 Learning in Unknown LDS
In this chapter we consider an easier task than control, namely that of learning in a dynamical system. The task of the learner is to predict the next observation from past observations alone, while being unable to influence the underlying dynamical system.
We show in the next few sections how the methodology of convex relaxations extends to prediction in dynamical systems. This chapter uses similar a similar parameterization as DACs in context of prediction, while leaving more intricate parameter-efficient convex relaxations to the next chapters.
@@ 11.1 Learning in dynamical systems
There are numerous different interpretations of “learning” in dynamical systems. For this chapter the notion we are concerned with is that of prediction. More precisely, consider a sequence of observable inputs to a dynamical system as in Figure 11.1. Our goal is to correctly predict the output sequence to minimize a given loss metric.
@@
For the rest of this chapter, we consider a partially observable linear dynamical system that is governed by the following equations.
@@
Here, as we have denoted throughout the text, $\mathbf{x}$ is the hidden state, $\mathbf{u}$ is the observed control input, $\mathbf{w},\mathbf{v}$ are perturbations/noises, and $\mathbf{y}$ is an observed linear transformation of the state. Throughout this chapter we assume that the system matrices $A,B,C$ are unknown to the learner.
@@ 11.2 Online learning of dynamical systems
In the problem of online learning in a linear dynamical system, sequentially for $t=1,...,T$, the online learning algorithm:
  • observes input $\mathbf{u}_{t-1}\in{\mathbb{R}}^{d_{u}}$
  • predicts output $\hat{\mathbf{y}}_{t}\in{\mathbb{R}}^{d_{y}}$
  • observes true system output $\mathbf{y}_{t}$
  • suffers loss according to a pre-specified loss function, $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, e.g.
@@
The goal of the online learner is to minimize total loss over all game iterations. If the observations ${\mathbf{y}}_{t}$ are arbitrary, this is an ill-defined goal, and rather than minimizing loss, we consider a comparative analogue.
More precisely, we consider a family of prediction rules from available information to prediction, and consider the metric of regret vs. the best of these rules in hindsight. An example of a rule could be a fixed linear map, i.e. $\pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=K\mathbf{u}_{t-1}$, where $K$ can potentially depend on the system matrices $A,B,C$.
Formally, let $\Pi$ be a family of mappings from $\{\mathbf{u}_{\tau},\mathbf{y}_{\tau}\}\mapsto\hat{\mathbf{y}}_{t}$, i.e.
$\pi\in\Pi\ ,\ \pi(\mathbf{u}_{1:t-1},\mathbf{y}_{1:t-1})=\hat{\mathbf{y}}_{t}^{\pi}\in{\mathbb{R}}^{d_{y}}.$
Let ${\mathcal{A}}$ be a prediction algorithm that iteratively produces predictions $\hat{\mathbf{y}}_{t}$. Then its regret w.r.t. $\Pi$ is defined as
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}.$
We proceed to define common prediction classes in the context of linear dynamical systems.
@@ 11.3 Classes of Prediction Rules
We now consider what would be a reasonable prediction rule for the output of a linear dynamical system? The case of full observation is particularly simple, and we deal with it first. We then consider prediction rules for partially observed linear dynamical systems.
However, before considering the prediction rules themselves, we discuss how to compare classes of such rules. This is analogous to the definitions and discussion in chapter 6 where we discuss comparison between policy classes for control. However, the definitions and conclusions for prediction are simpler.
@@
We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following holds. For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi\in\Pi_{1}$, there exists $\pi_{2}\in\Pi_{2}$ such that,
@@
The significance of this definition is that if a simple predictor class approximates another more complicated class, then it suffices to consider only the first.
The implication of $\varepsilon$-approximation on the regret is given below.
@@ 11.3.1 Linear Predictors
Perhaps the simplest prediction rule is linear, in either the observation, or input, or both. More precisely, the class of all predictions of the form
$\hat{\mathbf{y}}_{t+1}=\sum_{i=1}^{h}M_{1,i}\mathbf{y}_{y-i}+\sum_{j=1}^{k}M_{2,j}\mathbf{u}_{y-j}.$
We denote by $\Pi^{LIN}_{h,k,\kappa}$ the class of all linear predictors with $\sum_{i=0}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
@@
A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control. We denote by $\Pi_{h,k,\kappa}^{LIN}$ the set of all linear controllers bounded as $\sum_{i=0}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
11.3.2 Open Loop Optimal Predictors
An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open loop linear dynamical predictors, which we now define.
@@ 11.3.2 Open Loop Optimal Predictors
$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}.$
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next system output accurately. We denote by $\prod_{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
@@
where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{OLOP}$ the class of all such predictors.
@@
Suppose that $\|B\|+\|C\|\leq\gamma$, and $\|\mathbf{u}_{t}\|\leq U$ for all $t\in[T]$. Then the class $\prod_{h}^{OLOP}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod_{\star}$.
@@
Consider any sequence and any predictor from $\prod_{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$. Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}$ for the same system matrices. Then we have
@@
This lemma explains why the class $\prod^{OLOP}$ is an optimal prediction class for open loop prediction, in which the realized noise (through observations) cannot be used inside of the prediction rule, and in the case of full observation.
We note that the class $\prod^{LIN}$ also $\varepsilon$-approximates the class $\prod^{OLOP}$ for appropriate choice of parameters. This is left as an exercise (Exercise 11.6) to the reader.
@@ 11.4 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well some of the relationships between them. Notably, the class of open loop optimal predictors is optimal for certain scenarios and requires knowledge of only three matrices. However, the parametrization of this class is non-convex in the matrices $A,B,C$, which presents a computational challenge.
This challenge is circumvented by using the fact that the class $\prod^{LIN}$ $\varepsilon$-approximates the classe $\prod^{OLOP}$, and as we show in the next chapter, also optimal closed loop predictors. Further, the natural parametrization of $\prod^{LIN}$ is convex and thus can be learned by online convex optimization.
@@
Algorithm 6 Learning $\prod^{LIN}$ by Online Gradient Descent
@@
This algorithm is an instance of the online gradient descent algorithm, Theorem A.1 from Appendix A directly gives the following bound for it. We denote by $\mathcal{K}$ the set of all $h\times k$-tuple of matrices $M_{1,1:h},M_{2,1:k}$, such that they are bounded as needed. Then we obtain
@@
For choice of step sizes $\eta_{t}=O(\frac{1}{\sqrt{t}})$, and ${\mathcal{A}}$ being Algorithm 6, we have that
$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{LIN}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})=O(\sqrt{T}).$

### v4

@@ 8.3 Nonstochastic System Identification
The problem we study in this section is online nonstochastic system identification of a linear time-invariant dynamical system.
@@
Assuming that
@@
    the system is $(k,\kappa)$-strongly controllable,
@@
    the perturbations $\mathbf{w}_{t}$ do not depend on $\mathbf{u}_{t}$,
and for $T_{0}=\Omega\left(\frac{d^{2}k^{2}\kappa^{2}\gamma^{6}}{\varepsilon^{2}}\log\frac{1}{\Delta}\right)$, the following holds. Algorithm 5 outputs a pair $(\widehat{A},\widehat{B})$ that satisfies, with probability $1-\Delta$,
@@
First note that for any $t$, since the system is $\gamma$ stable and the controls have norm $\sqrt{d}$, the state $\mathbf{x}_{t}$ is bounded as
$\|\mathbf{x}_{t}\|=\left\|\sum_{i=1}^{t}A^{i-1}\left(B\eta_{t-i}+\mathbf{w}_{t-i}\right)\right\|\leq 2\sqrt{d}\gamma.$
@@
By this display, we already have $\|B-\widehat{B}\|\leq\varepsilon$. We will make use of this observation to establish the proximity of $\widehat{A}$ and $A$, whose proof is left as Exercise 8.7.
Note that $\widehat{A}$ does not necessarily satisfy $\widehat{A}\widehat{C_{0}}=\widehat{C_{1}}$. Hence we require the following lemma.
@@
We now use this lemma on the normal equations, i.e. $AC_{0}C_{0}^{\top}=C_{1}C_{0}^{\top}$ and $\widehat{A}\widehat{C_{0}}\widehat{C_{0}}^{\top}=\widehat{C_{1}}\widehat{C_{0}}^{\top}$. First, observe that for any $\varepsilon<1$
@@ 8.4 From Indentification to Nonstochastic Control
Having recovered approximate estimates $(\widehat{A},\widehat{B})$ of the system matrices, a natural question arises: how well does the Gradient Perturbation Algorithm 4 perform on the original system $(A,B)$ when provided with only approximate estimates of the same? We briefly state a result to this extent, whose proof we leave to the explorations of an interested reader.
@@
This result translates to a regret bound of $\tilde{O}(T^{2/3})$ for nonstochastic control of unknown, yet stable, linear systems, by first running the nonstochastic system identification procedure, and subsequently using the estimates thus obtained to run GPC. Details of how to derive this regret bound are left as Exercise 8.7.
8.5 Summary and Key Takeaways
This chapter explored online nonstochastic system identification, where a controller must simultaneously learn an unknown system model while making control decisions in an adversarial setting. Unlike classical stochastic system identification, which assumes structured noise and statistical convergence, the nonstochastic approach makes no such assumptions and instead relies on regret minimization techniques to ensure efficient learning.
The key insights from this chapter are as follows.
  • The Exploration-Exploitation Tradeoff A fundamental challenge in online system identification is balancing exploration, which enables learning the system dynamics, and exploitation, which minimizes control cost. Too much exploration leads to high immediate costs, while too much exploitation can result in poor long-term learning and suboptimal control performance. The algorithms in this chapter address this by incorporating controlled perturbations in the control inputs, ensuring sufficient exploration while keeping regret low. However, the tradeoff results in a regret of $O(T^{2/3})$, rather than $O(\sqrt{T})$ we had in the last chapter.
    More sophisticated techniques for balancing exploration and exploitation can result in better regret bounds, some of which are surveyed in the bibliographic section.
  • Regret Minimization as a Learning Metric Unlike traditional identification methods that aim for asymptotic convergence, online nonstochastic identification is evaluated in terms of regret, measuring performance relative to the best fixed model in hindsight.
8.5.1 Looking Ahead
This part of the book considered the online nonstochastic control problem, the regret metric, and online learning techniques applied to control. We have thus far considered the case where the state is observable to the controller, an assumption which doesn’t always hold.
In the next parts of the book, we consider the setting of partial observability, where the system state is unknown, but an observed signal is available. We start with the problem of learning, and move onto online control.
@@ 8.7 Exercises
8.2. Design an analogous control procedure as that for Exercise 8.7 with an identical regret upper bound (up to constants), when the horizon $T$ is not known in advance.
@@
Chapter 9 Learning in Unknown Linear Dynamical Systems
In previous chapters, we focused on control in dynamical systems, often assuming that the system state was fully observable and that system dynamics were either known or could be estimated through system identification. In many real-world scenarios, however, these assumptions do not hold. Instead, an agent may only have access to partial observations of the system and must learn to make accurate predictions without directly influencing the system’s evolution.
This chapter focuses on learning in a dynamical system, which means predicting future system outputs based on past observations while the underlying system parameters remain unknown. Unlike control, where an agent actively influences system trajectories to minimize a given cost, learning in this setting is a passive task—the goal is to accurately forecast future outputs without modifying the system’s behavior.
The ability to predict future states and outputs of an unknown system is fundamental to many domains, including time-series forecasting, financial modeling, sensor networks, autonomous systems, and most recently large language models. Moreover, learning a predictive model is often a key stepping stone toward solving more complex problems, such as control under partial observability or reinforcement learning in dynamical systems. In this chapter, we study online learning algorithms for prediction in unknown linear dynamical systems, with an emphasis on regret minimization.
@@ 9.1 Learning in Dynamical Systems
The term learning in dynamical systems can refer to a variety of tasks, including state estimation, system identification, and policy learning. In this chapter, we focus on the specific task of prediction, where the objective is to estimate future outputs of a dynamical system based on past observations. Unlike classical system identification, which seeks to recover the system parameters explicitly, the learning approach taken here aims to optimize predictions directly.
To formally define the problem, consider a sequence of observable inputs fed into a dynamical system, as illustrated in Figure 9.1. The learner observes these inputs and aims to predict the corresponding outputs while minimizing a given loss function.
@@
For the remainder of this chapter, we consider a partially observable linear time invariant dynamical system described by the following equations. The extention of the material in this chapter to time varying linear dynamical systems is left as exercise 9.7.
@@
Here, $\mathbf{x}_{t}$ represents the hidden system state, $\mathbf{u}_{t}$ is the observed control input, and $\mathbf{y}_{t}$ is the measured system output. The terms $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ account for process and observation noise, respectively. A crucial challenge in this setting is that the system matrices $A,B,C$ are unknown to the learner.
The learner’s objective is to construct an efficient prediction rule that estimates $\mathbf{y}_{t}$ based on past observations, despite the unknown system parameters.
@@ 9.2 Online Learning of Dynamical Systems
In an online learning setting, the learner sequentially observes data and makes predictions in real time. Formally, the process unfolds as follows:
  1. 1.
    At each time step $t$, the learner observes the control input $\mathbf{u}_{t-1}$.
  2. 2.
    The learner produces a prediction $\hat{\mathbf{y}}_{t}$ for the next output.
  3. 3.
    The true output $\mathbf{y}_{t}$ is revealed.
  4. 4.
    The learner incurs a loss based on the difference between $\mathbf{y}_{t}$ and $\hat{\mathbf{y}}_{t}$, using a predefined loss function $\ell(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$, such as squared error:
@@
The goal is to minimize the total loss over time. However, in adversarial or unpredictable settings, direct loss minimization may be ill-posed. Instead, we consider a regret-based formulation, where the learner’s performance is measured relative to the best possible predictor from a predefined class of functions.
More precisely, let $\Pi$ be a family of prediction rules mapping past observations to predictions. The regret of an online learning algorithm $A$ with respect to $\Pi$ is given by:
$\text{Regret}_{T}(A,\Pi)=\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|^{2}-\min_{\pi\in\Pi}\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}^{\pi}_{t}\|^{2}$
where $\hat{\mathbf{y}}^{\pi}_{t}$ denotes the prediction made at time $t$ by the best function $\pi\in\Pi$ in hindsight.
This regret-based framework allows us to compare different classes of predictors and design learning algorithms that adapt efficiently to unknown dynamical systems.
@@ 9.3 Classes of Prediction Rules
We now explore what constitutes a reasonable prediction rule for the output of a linear dynamical system. However, before considering the prediction rules themselves, we discuss how to compare classes of such rules. This is analogous to the definitions and discussion in chapter 6 where we discuss comparison between policy classes for control. However, the definitions and conclusions for prediction are simpler.
@@
We say that a class of predictors $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following holds. For every sequence of $T$ inputs and outputs $\{\mathbf{u}_{1:T},\mathbf{y}_{1:T}\}$, and every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that,
@@
The significance of this definition is that if a simple predictor class approximates another more complicated class, then it suffices to consider only the first. The implication of $\varepsilon$-approximation on the regret is given below.
@@ 9.3.1 Linear Predictors
The simplest prediction rule is a linear function of the observations, inputs, or both. More precisely,
@@
A linear controller parametrized by matrices $\{M_{1,i},M_{2,j}\}\in{\mathbb{R}}^{h\times d_{y}\times d_{y}+k\times d_{y}\times d_{u}}$ is a linear operator mapping state to control as
$\hat{\mathbf{y}}_{t}=\sum_{i=1}^{h}M_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{k}M_{2,j}\mathbf{u}_{t-j}.$
We denote by $\Pi_{h,k,\kappa}^{L}$ the set of all linear controllers bounded as $\sum_{i=1}^{h}\|M_{1,i}\|+\sum_{j=1}^{k}\|M_{2,j}\|\leq\kappa$.
9.3.2 Linear Dynamical Predictors
An intuitive class of prediction rules that makes use of the special structure of linear dynamical systems is that of open-loop linear dynamical predictors, which we now define.
@@ 9.3.2 Linear Dynamical Predictors
$=\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t}\mathbf{x}_{0}.$
Thus, the operator $\pi_{\{A,B,C,\mathbf{x}_{0}\}}$ takes a history of states and controls and gives the next system output accurately. We denote by $\prod^{\star}$ the set of all such predictors, that in the case of zero noise give a perfectly accurate prediction.
@@
where we define all $\mathbf{u}$ vectors in negative times as zero. Denote by $\prod_{h}^{\star}$ the class of all such predictors.
@@
Suppose that $\|B\|+\|C\|\leq\gamma$, and $\|\mathbf{u}_{t}\|\leq U$ for all $t\in[T]$. Then the class $\prod_{h}^{\star}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod^{\star}$.
@@
Consider any sequence and any predictor from $\prod^{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$. Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}^{\star}$ for the same system matrices. Then we have
@@ 9.4 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well some of the relationships between them. Notably, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices. However, the parametrization of this class is non-convex in the matrices $A,B,C$, which presents a computational challenge.
This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$. Further, the natural parametrization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
@@
Algorithm 6 Learning $\prod^{L}$ by Online Gradient Descent
@@
This algorithm is an instance of the Online Gradient Descent algorithm, and Theorem 6.1 directly gives the following bound for it. We denote by $D,G$ the diameter of the set $\mathcal{K}$ as in Algorithm 6, and by $G$ an upper bound on the norm of the gradients of the loss functions $\ell_{t}$.
@@
For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, and ${\mathcal{A}}$ being Algorithm 6, we have that
$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{L}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})\leq 2GD\sqrt{T}.$
9.5 Summary
In this chapter, we explored the problem of learning in unknown linear dynamical systems, where the goal is to predict future system outputs based on past observations while the underlying system parameters remain unknown. Unlike traditional control settings where an agent actively influences system behavior, the learning problem here is passive, focusing on accurate forecasting without intervention.
We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations. We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric. The regret framework allows for direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
A key takeaway from this chapter is that learning in LDS requires balancing expressivity and computational efficiency. Simple linear predictors may be computationally efficient but lack expressivity, whereas more sophisticated dynamical predictors require careful algorithmic design to remain tractable. The regret analysis provided insights into how online learning algorithms adapt to unknown system parameters while ensuring competitive performance.
This discussion sets the stage for the next chapter on Kalman filtering, where we transition from open-loop predictors to closed-loop estimation techniques that integrate feedback from observations to refine state estimates.